# Magneto-optical response of Dynes superconductors

### v1

@@ 3.1 Voigt geometry
Figure 3: Density of states (DOS) in the vicinity of the Fermi energy normalized to the DOS in the normal state for magnetic fields up to 2.8 T. The DOS of the HH and the SBW models are slightly different even for B=0 due to the slight difference in the zero field fits. The DOS was calculated with the Dynes formula, Eq. 3, using parameters from the fit by the HH model (solid lines), and with the SBW model using parameters from the fit of the THz data (dashed lines).
First, the SBW model was used to fit the zero-field experimental data using the zero-field gap $\Delta_{0}$, the DC conductivity $\sigma_{0}$ and the pair-breaking parameter $\alpha$ as free parameters. In the magnetic field, $\alpha(B)$ is the only free parameter of the fit since other variables are not field-dependent. The field dependences of the pair-correlation gap $\Delta(B)$ and of the spectroscopic gap $\Omega_{G}(B)$ were determined from the values of $\alpha(B)$ within the SBW model [5]. Measurements up to 3.6 T revealed the suppression of the pair-correlation gap $\Delta(B)$ and of the spectroscopic gap $\Omega_{G}(B)$. Above 2.4 T, $\Omega_{G}=0$ but $\Delta(B)\neq 0$ and the superconductor enters a gapless state until the superconductivity is completely suppressed at 3.5 T. The SBW model correctly predicts a shift in $\Omega_{G}$ with the applied magnetic field, but it fails to describe the shape of the real part of the THz conductivity $\sigma_{1}(\omega)$ in detail, see solid lines in Fig 1.
We decided to compare the results of the SBW model with the theoretical prediction of Herman and Hlubina [2]. First, we fitted the experimental data in zero magnetic field with the DC conductivity $\sigma_{0}$, the zero-field gap $\Delta_{\Gamma\rightarrow 0}$, the pair-conserving scattering rate $\Gamma_{s}$ and the pair-breaking scattering rate $\Gamma$ as free parameters. The value of the gap $\Delta$ is is approximated by the formula  [24, 25]
@@
which is valid in the limit of zero temperature. For non-zero magnetic fields, we used the resulting values except for the pair-breaking scattering rate $\Gamma(B)$ which was treated as a free parameter. The magnetic field dependence of gap $\Delta(B)$ is governed by $\Gamma(B)$ using equation 9. The conductivity $\sigma(\nu)$ predicted by the HH model, see dashed lines in Fig 1, decreases more slowly at low frequencies. For low magnetic fields, the SBW model seems to be more accurate, but from 1 T, the HH model describes more precisely the experimental results. In Fig. 2, we compare the values of the parameters of the SBW and of the HH theoretical models. The pair-correlation gap $\Delta$ and $\Gamma$, respectively $\alpha$, have qualitatively and quantitatively the same dependence on the magnetic field. To better understand the difference between these models, we plot the density of states within the SBW model and compare it to the Dynes formula justified by the HH model. The parameters for the DOS are taken from the fits of the complex conductivity $\sigma(\nu)$ with each model. They are qualitatively different. While the SBW DOS has no states for $|E|<\Omega_{G}$ for $B<2.4$ T, the Dynes formula fills the gap gradually, see Fig 3.
3.2 Faraday geometry
We measured the optical response of three NbN films of different thicknesses (5.3, 11.5, and 30.1 nm) using time-domain THz spectroscopy. The experimental setup, as well as the treatment used to extract the THz conductivity spectra from our data, are reported in detail in [9].
In the Faraday geometry, the magnetic field and consequently the vortices are perpendicular to the plane of the film. Their presence substantially alters the zero-field properties in type II superconductors. To interpret the experimental results, we use a model in which the vortices are approximated as normal-state cylinders surrounded by a superconducting matrix. The spectral response of the superconducting matrix is provided by the HH model [2] where $\Gamma$ changes with the magnetic field and describes the pair-breaking effect of the field. The gap is modified due to $\Gamma$ according the equation 9.
Table 1: Parameters of NbN films: $d_{f}$ film thickness, $T_{c}$ critical temperature, $\Gamma_{s}$ field independent value of pair-conserving scattering rate, $\Delta_{\Gamma\rightarrow 0}$ value of the gap corresponding to $\Gamma\rightarrow 0$ and $\sigma_{0}$ DC conductivity
@@ 3.2 Faraday geometry
Figure 4: Real (a, c, e) and imaginary (b, d, f) parts of the THz conductivity of NbN thin films measured in the Faraday geometry at 2 K up to 7 T. Three NbN samples of different thicknesses were studied: (a-b) $d_{f}=5.3$ nm, (c-d) $d_{f}=11.5$ nm and (e-f) $d_{f}=30.1$ nm. Symbols: experimental data, lines: fits using the Maxwell-Garnett theory and the HH model.
In the THz range, the conditions for the so-called long-wavelength limit are fulfilled: the wavelengths are much longer than the distances between the vortices, thus the inhomogeneous system can be represented by an effective optical conductivity. The geometry of the normal state inclusions inside the superconducting matrix can be well described by the Maxwell-Garnett model [10] which yields the effective conductivity ${\tilde{\sigma}}_{\rm MG}$:
@@
where $f_{\rm n}$ and ${\tilde{\sigma}}_{\rm n}$ are the volume fraction and the conductivity of the normal state vortex cores, respectively; further, ${\tilde{\sigma}}_{\rm s}$ the conductivity of the superconducting matrix. $K$ is the shape factor of inclusions and has been taken as $K=1$ which corresponds to the value for cylinders with their axis perpendicular to the electric field.
The THz conductivity spectra of the three NbN films are shown in Fig. 4. In our fits, we used two free parameters: the volume fraction of the vortex cores $f_{\rm n}$, and the pair-breaking scattering rate $\Gamma$ which influences the shape of the superconducting conductivity spectra $\sigma_{s}(\nu)$. The remaining parameters are taken from the zero field fit and they are field independent. In Fig. 4 (e) and (f), we show illustrative examples of fits using the zero-field $\sigma_{s}(\nu)$ (dashed curves) in order to demonstrate the importance of $\Gamma(B)$.
The pair-breaking scattering rate $\Gamma$ is proportional to the magnetic field for all 3 studied NbN films, see Fig. 5(b). The gap $\Delta$ (Fig. 5(c)) was evaluated using equation 9 from the fitted values of $\Gamma$.
The normal fraction should be a linear function of the magnetic field ($f_{n}=B/B_{c2}$), but this holds only for the thickest sample, see Fig. 5(a). The square root dependence $f_{n}\propto aB^{1/2}$ observed by Xi et al. [11] successfully describes $f_{n}(B)$ for the 5 nm and the 11.5 nm thick NbN films. A similar square root dependence was also observed in overdoped La2-xCexCuO4 [31].
The sublinear behaviour can be attributed to the limitations of the simple effective medium model or possibly due to the shrinking of the vortex cores [32, 33] with increasing magnetic field. Finally, the upturn in the real part of the conductivity $\sigma_{1}(\omega)$ at low frequencies which is not fully described by our model results from the vortex dynamics [12], since this effect can not taken into account in the effective medium approaches.
@@
4 Conclusions
The theoretical model by Herman and Hlubina [2] for Dynes superconductors allows us to calculate the optical conductivity for varying ratios of the pair-conserving and pair-breaking processes. We have demonstrated that the HH model can well describe the optical conductivity in magnetic fields both in Voigt and Faraday geometries by varying the pair-breaking scattering rate $\Gamma$. In this way, the model serves as an effective phenomenology describing the pair-breaking effect of the increasing external magnetic field suppressing the $T_{c}$, filling the gap in the density of states, and affecting the optical conductivity. In the Voigt geometry, the HH model offers a better theoretical description of the optical conductivity of the Nb thin film measured by Lee et al. [1] than the SBW model used earlier. In the Faraday geometry, the THz response of the vortex lattice is described by the Maxwell-Garnett effective medium theory [10] with the properties of the superconducting matrix following the HH model [2]. As the range of the found parameters $\Gamma\lesssim\Delta\lesssim\Gamma_{s}/10$ fulfills the natural expectations and suggests possible usage of the HH model in the dirty limit [2], the presented formulation allows for the explicit presence of the pair-conserving scattering. This may be useful in the analysis of cleaner systems.
Last but not least, let us emphasize that the presented analysis utilizes the HH model as an effective phenomenology with the fewest free parameters, taking into account $\Gamma(B)$ as the effective pair-breaking scattering rate. This approach allows for straightforward comparison with the related results reported in the current literature. The difference in shapes of $\Gamma(B)$ dependence considering different geometries originates probably from the combination of the Zeeman splitting and spin-orbit interaction in Voigt and from the orbital effects in the Faraday geometry. To proceed further with the analysis of the presented experiments and distinguish the role of individual effects on terahertz conductivity measurements in more detail, one would need a more elaborate description of the Zeeman splitting and of the spin-orbit interaction, to take into account the screening current generation (orbital effects), as well as the role of intrinsic pair-breaking and pair-conserving disorder. In the best-case scenario, the model could go beyond the Born approximation limit, allowing for higher magnitudes of the individual effects.
@@ Acknowledgements
We are grateful to K. Ilin and M. Siegel for preparing and characterizing the NbN sample, as well as to Jae Hoon Kim and Ji Eun Lee, who were very kind to provide their data [1]. We also acknowledge the financial support by the Czech Science Foundation (Project No. 21-11089S), by European Union and the Czech Ministry of Education, Youth and Sports (Project TERAFIT - CZ.02.01.01/00/22 008/0004594) and by the INTER-COST (Project LUC24098). This work was also supported by the Slovak Research and Development Agency under the Contract no. APVV-23-0515 and by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 945478.

### v3

@@ III.1 Voigt geometry
Figure 5: Density of states (DOS) in the vicinity of the Fermi energy normalized to the DOS in the normal state for magnetic fields up to 2.8 T. The DOS of the HH and the SBW models are slightly different even for B=0 due to the slight difference in the zero field fits. The DOS was calculated with the Dynes formula, Eq. (3), using parameters from the fit by the HH model (solid lines), and with the SBW model using parameters from the fit of the THz data (dashed lines).
First, the SBW model was used to fit the zero-field experimental data using the zero-field gap $\Delta_{0}$, the DC conductivity $\sigma_{0}$, and the pair-breaking parameter $\alpha$ as free parameters. In magnetic field, $\alpha(B)$ is the only free parameter of the fit since other variables are not field-dependent. The field dependences of the pair-correlation gap $\Delta(B)$ and of the spectroscopic gap $\Omega_{G}(B)$ were determined from the values of $\alpha(B)$ within the SBW model [4]. Measurements from zero to 3.6 T revealed the suppression of the pair-correlation gap $\Delta(B)$ and of the spectroscopic gap $\Omega_{G}(B)$. Above 2.4 T, $\Omega_{G}=0$ but $\Delta(B)\neq 0$ and the superconductor enters a gapless state until the superconductivity is completely suppressed at 3.5 T. The SBW model correctly predicts a shift in $\Omega_{G}$ with the applied magnetic field, but it does not fully describe the shape of the real part of the THz conductivity $\sigma_{1}(\omega)$ at lower frequencies, see solid lines in Fig 3. We compared the results of the SBW model with the theoretical prediction of Herman and Hlubina [8]. First, we fitted the experimental data in a zero magnetic field with DC conductivity $\sigma_{0}$, the zero-field gap $\Delta_{\Gamma\rightarrow 0}$, the pair-conserving scattering rate $\Gamma_{s}$, and the pair-breaking scattering rate $\Gamma$ as free parameters. The value of the gap $\Delta$ was approximated by the formula  [26, 29]
@@
which is valid in the limit of zero temperature. For non-zero magnetic fields, we used the resulting values except for the pair-breaking scattering rate $\Gamma(B)$ which was treated as a free parameter. The magnetic-field dependence of the gap $\Delta(B)$ is governed by $\Gamma(B)$ via Eq. (9). The real part of the conductivity $\sigma_{1}(\nu)$ predicted by the HH model, see dashed lines in Fig. 3, decreases more slowly at low frequencies due to the presence of the sub-gap absorption [8]. For low magnetic fields, the SBW model seems to be more accurate, but from 1 T, the HH model describes the experimental results better. In Fig. 4, we compare the values of the parameters of the SBW and of the HH theoretical models. The pair-correlation gap $\Delta$ and $\Gamma$, as well as $\alpha$, have qualitatively and quantitatively the same magnetic field dependence. To better understand the difference between these models, we plot the density of states within the SBW model and compare it to the Dynes formula justified by the HH model. The parameters for the DOS are taken from the fits of the complex conductivity $\sigma(\nu)$ with each model. They are qualitatively different. Whereas the SBW DOS has no states for $|E|<\Omega_{G}$ and $B<2.4$ T, the DOS calculated with the Dynes formula using the parameters obtained from the fit by the HH model fills the gap gradually, see Fig. 5.
III.2 Faraday geometry
We measured the optical response of three NbN films of different thicknesses (5.3, 11.5, and 30.1 nm) using time-domain THz spectroscopy. In brief, the samples were placed in a He-bath optical cryostat (Spectromag, Oxford Inst.) comprising a superconducting magnet, and their complex THz transmittance was measured using broadband THz pulses. A dedicated procedure was applied to determine the conductivity spectra of the films, whose thicknesses are about five orders of magnitude lower than those of the substrates. Further details of the experimental setup, as well as of the treatment used to extract the THz conductivity spectra from our data, are reported in detail in [9].
In the Faraday geometry, the magnetic field is perpendicular to the plane of the film, and consequently, so are the vortices. Their presence substantially alters the zero-field properties in type-II superconductors. To interpret the experimental results, we use a model in which the vortices are approximated by normal-state cylinders surrounded by a superconducting matrix. For fields reported here ($\mu_{0}H>1$ T), the distance between vortices is smaller than the penetration depth $\lambda_{L}$, so the magnetic flux penetrates into the superconducting matrix. The spectral response of the superconducting matrix is described by the HH model [8] which incorporates the pair-breaking effects of the magnetic field through the parameter $\Gamma(B)$. Consequently, the gap is modified due to $\Gamma$ according to Eq. (9).
Table 1: Parameters of the measured NbN films—$d_{f}$: film thickness, $T_{c}$: critical temperature, $\Gamma_{s}$: field-independent value of pair-conserving scattering rate, $\Delta_{\Gamma\rightarrow 0}$: energy of the gap corresponding to $\Gamma\rightarrow 0$, and $\sigma_{0}$: DC conductivity
@@ III.2 Faraday geometry
Figure 6: Real (a, c, e) and imaginary (b, d, f) parts of the THz conductivity of NbN thin films measured in the Faraday geometry at 2 K up to 7 T. Three NbN samples of different thicknesses were studied: (a-b) $d_{f}=5.3$ nm, (c-d) $d_{f}=11.5$ nm and (e-f) $d_{f}=30.1$ nm. Symbols: experimental data; full lines: fits using the Maxwell-Garnett theory and the HH model; dashed lines: fits neglecting $\Gamma(B)$.
In the THz range, the conditions for the so-called long-wavelength limit are fulfilled: the wavelengths are much longer than the distances between the vortices, thus the inhomogeneous system can be represented by an effective optical conductivity. The geometry of the normal-state inclusions within the superconducting matrix can be well described within the Maxwell-Garnett model [10] which yields the effective conductivity ${\tilde{\sigma}}_{\rm MG}$:
@@
where $f_{\rm n}$ and ${\tilde{\sigma}}_{\rm n}$ are the volume fraction and the conductivity of the normal-state vortex cores, respectively, and ${\tilde{\sigma}}_{\rm s}$ is the conductivity of the superconducting matrix. $K$ is the shape factor of inclusions which was set to $K=1$, corresponding to the value for cylinders with their axes perpendicular to the electric field, as in the present case.
The THz conductivity spectra of the three NbN films are shown in Fig. 6. In our fits, we used two free parameters: the volume fraction of the vortex cores $f_{\rm n}$, and the pair-breaking scattering rate $\Gamma$ which influences the shape of the superconducting conductivity spectra $\sigma_{s}(\nu)$. The remaining parameters are field independent, and their values were taken from the zero-field fit. Fits in the Figure 6(e, f) compare the following calculations within the Maxwell-Garnett theory (MGT). Both use HH model for $\sigma_{\mathrm{s}}(\nu)$. However, dashed curves use zero magnetic field $\Gamma(0)$, while solid curves account for magnetic-field dependence through $\Gamma(B)$. The latter demonstrates improved agreement with experimental data. The pair-breaking scattering rate $\Gamma(B)$ increases with the magnetic field for all three NbN films, as shown in Fig. 7(b). The gap energy $\Delta$ [shown in Fig. 7(c)] was evaluated using Eq. (9) from the fitted values of $\Gamma$.
The normal fraction should be a linear function of the magnetic field ($f_{n}=B/B_{c2}$), but this holds only for the thickest sample, see Fig. 7(a). The square root dependence $f_{n}\propto aB^{1/2}$ observed by Xi et al. [11] successfully describes $f_{n}(B)$ for the 5 nm and the 11.5 nm thick NbN films. A similar square-root dependence was also observed in overdoped La2-xCexCuO4 [35]. The sublinear behaviour can be attributed to the limitations of the simple effective medium model, or possibly to the shrinking of the vortex cores [36, 37] with increasing magnetic field.
It is important to note that the presented theoretical description is incomplete without accounting for vortex motion in the Faraday geometry. Presently, no unified model exists that would simultaneously incorporate (i) the BCS theory description of optical conductivity, (ii) the presence of vortex cores, and (iii) the vortex-motion-induced absorption. In particular, the THz regime exhibits a distinct behavior compared to the GHz range: on picosecond timescales, the vortex motion is limited to localized oscillations around pinning sites, and the effective potential barriers may differ from those governing the low-frequency dynamics.
The contribution of vortices to conductivity can be described by the formula presented in the review paper by Pompeo et al. [38] which takes the same form for the Coffey-Clem [13] and Brandt [14] models. However, the models differ in their physical interpretations. A key parameter is the depinning frequency $\nu_{p}$, above which the driving force dominates the pinning, leading to dissipative flux flow. An estimate from measurements by Basistha et al. [39] suggests that the depinning frequency of NbN is in the MHz range, while for niobium, Pompeo et al. [40] report $v_{p}=7.2$ GHz. This confirms that THz waves act deep within the flux-flow regime, where vortices are highly mobile.
In our prior work [12], we analysed vortex dynamics in the NbN film deposited on Si substrate (the one examined also here) and demonstrated the role of the vortex motion in the upturn of $\sigma_{1}$ at low THz frequencies which is not accounted by the present model. This underscores the necessity of developing an integral description of the vortex dynamics to fully capture the electrodynamic response of superconductors.
@@
IV Conclusions
The theoretical model by Herman and Hlubina [8] for conventional superconductors allowed us to calculate the optical conductivity for varying ratios of the pair-conserving and pair-breaking processes. We demonstrated that the HH model can well describe the optical conductivity in magnetic fields both in Voigt and Faraday geometries owing to the varying pair-breaking scattering rate $\Gamma$. In this way, the model serves as an effective phenomenology describing the pair-breaking effect of the increasing external magnetic field which suppresses the $T_{c}$, fills the gap in the density of states, and affects the optical conductivity. In the Voigt geometry, the HH model offers a better theoretical description of the THz conductivity of the Nb thin film measured by Lee et al. [5] than the SBW model used previously. In contrast with the SBW model, it predicts the existence of a non-zero DOS at the Fermi energy even for $B=0$ T, as well as the filling of the gap region with an increasing magnetic field. Subgap absorption described by the gapless regime in the HH model better describes measurements of $\sigma_{1}(\nu)$ at low frequencies. Comparison of the SBW model and HH model based analysis therefore opens an interesting question about the presence of the gapped-superconducting state once $B\neq 0$.
In the Faraday geometry, the THz response of the vortex lattice is described by the Maxwell-Garnett effective medium theory [10] assuming a superconducting matrix following the HH model [8]. As one would expect, the values of resulting parameters fulfill relations $\Gamma\lesssim\Delta\lesssim\Gamma_{s}/10$, suggesting the possibility of applying the HH model in the dirty limit [8]. However, the presented formulation allows for an explicit presence of the pair-conserving scattering; this may be useful for analyzing cleaner superconducting systems.
Last but not least, let us emphasize that the presented analysis utilizes the HH model as an effective phenomenology with the fewest free parameters, taking into account $\Gamma(B)$ as the effective pair-breaking scattering rate. This approach allows for a straightforward comparison with related results reported recently in the literature. The difference in shapes of $\Gamma(B)$ dependences considering different geometries probably originates from a combination of the Zeeman splitting and the spin-orbit interaction in the Voigt geometry and from the orbital effects in the Faraday geometry. To proceed further with the analysis of the presented experimental data, and to better distinguish the consequences of individual effects on the THz conductivity measurements, one would need a more elaborate description of the Zeeman splitting and of the spin-orbit interaction, to take into account the screening current generation (orbital effects), as well as the role of intrinsic pair-breaking and pair-conserving disorder. In the best-case scenario, such an improved model could go beyond the Born approximation limit, allowing for higher magnitudes of the individual effects.
@@ Acknowledgements
We are grateful to K. Ilin and M. Siegel for preparing and characterizing the NbN sample, as well as to J. H. Kim and J. E. Lee who were very kind to provide their data [5]. We are also thankful to R. Hlubina and L. Skrbek for discussions. We acknowledge the financial support from the Czech Science Foundation (Project No. 21-11089S), by the European Union and the Czech Ministry of Education, Youth and Sports (Project TERAFIT - CZ.02.01.01/00/22_008/0004594) and by the INTER-COST (Project LUC24098). This work was also supported by the Slovak Research and Development Agency under the Contract no. APVV-23-0515 and by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 945478.